OpenStudy (anonymous):

solve for x: x^4-14x^2+13=0

OpenStudy (anonymous):

\[x^4-14x^2+13=0\]

OpenStudy (anonymous):

do you know how to solve this equation? \[y^2-14y+13 = 0\]

OpenStudy (anonymous):

mmmmm not really is cuz they gave us a reviw nd i dont remember anything from last year

OpenStudy (anonymous):

starting with your original problem, lets substitute y = x^2 into it. then we get the equation i posted above. To solve, you want to think of two numbers that multiply to 13, but add to make -14. Those two numbers are -13 and -1. so we have: \[y^2-14y+13 = (y-13)(y-1) = 0 \iff y = 13 , y = 1\] but now lets put back our substitution: \[y = 13 \iff x^2 = 13 \iff x = \pm \sqrt{13}\] and \[y = 1 \iff x^2 = 1 \iff x = \pm 1\]

OpenStudy (anonymous):

ok thanks a lot =)